Calculate the temperature to which a .400 L sample of gas at 1.00 atm and 0 degrees celcius must be heated at a constant pressure for the volume to change to 1.00L
1
Expert's answer
2013-07-12T09:21:50-0400
If P = const we can use Charles's law. Due to it we have
V1/T1 = V2/T2
From this equation we get that
T2 = (T1*V2)/V1 = (273 K*1 L)/4 L = 68.25K = -204.75 degrees celsius.